OpenStudy (anonymous):

Which equation represents the problem? The circumference of a circle equals the product of two, pi, and the radius. Use C for circumference, 3.14 for pi, and r for the radius.

OpenStudy (anonymous):

C=2*3.14*r

OpenStudy (anonymous):

Thanks. Here is another one. Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side?

OpenStudy (anonymous):

apply pathagorous therom

OpenStudy (anonymous):



OpenStudy (anonymous):

15^2=225 8^2=64 2x

OpenStudy (anonymous):

Right?